Senior Life Co. is an HMO for businesses in the Portland area. The following account balances appear on the balance sh...

  1. Senior Life Co. is an HMO for businesses in the Portland area. The following account balances appear on the balance sheet of Senior Life Co.: The stock outstanding when a corporation has issued only one class of stock.Common stock (330,000 shares authorized; 5,000 shares issued), $125 A dollar amount assigned to each share of stock.par, $625,000; Capital contributed to a corporation by the stockholders and others.Paid-In Capital in excess of par— common stock, $125,000; and Net income retained in a corporation that is reported as part of stockholders' equity.Retained earnings, $7,500,000. The board of directors declared a 1% stock dividend when the market price of the Shares of ownership of a corporation.stock was $159 a share. Senior Life Co. reported no income or loss for the current year.

    If an amount box does not require an entry, leave it blank. If no entry is required, select "No entry required" from the dropdown.

    a1. Journalize the entry to record the declaration of the Distributions of a corporation's earnings to stockholders.dividend, capitalizing an amount equal to market value.

    • Cash
    • No Entry Required
    • Paid-In Capital in Excess of Par-Common Stock
    • Stock Dividends Distributable
    • Stock Dividends
    • Cash
    • No Entry Required
    • Stock Dividends Distributable
    • Stock Dividends
    • Treasury Common Stock
    • Cash
    • Common Stock
    • No Entry Required
    • Paid-In Capital in Excess of Par-Common Stock
    • Stock Dividends

    Feedback

    a2. Journalize the entry to record the issuance of the stock certificates.

    • Common Stock
    • No Entry Required
    • Paid-In Capital in Excess of Par-Common Stock
    • Stock Dividends
    • Stock Dividends Distributable
    • Common Stock
    • No Entry Required
    • Paid-In Capital in Excess of Par-Common Stock
    • Stock Dividends Distributable
    • Stock Dividends

    Feedback

    b. Determine the following amounts before the stock dividend was declared: (1) total paid-in capital, (2) total retained earnings, and (3) total The owners' equity in a corporation.stockholders' equity.

    Total paid-in capital $
    Total retained earnings $
    Total stockholders' equity $

    c. Determine the following amounts after the stock dividend was declared and closing entries were recorded at the end of the year: (1) total paid-in capital, (2) total retained earnings, and (3) total stockholders' equity.

    Total paid-in capital $
    Total retained earnings $
    Total stockholders' equity $
Answer #1
a1
Stock Dividends 7950 =5000*1%*159
      Stock Dividends Distributable 6250 =5000*1%*125
      Paid-In Capital in Excess of Par-Common Stock 1700
a2
Stock Dividends Distributable 6250
      Common Stock 6250
b
Total paid-in capital 750000 =625000+125000
Total retained earnings 7500000
Total stockholders' equity 8250000
c
Total paid-in capital 757950 =625000+125000+7950
Total retained earnings 7492050 =7500000-7950
Total stockholders' equity 8250000